Current assets $ 11,000 Current liabilities $ 12,000
Noncurrent assets 88,000 Noncurrent liabilities 56,000
Stockholders’ equity 31,000

The company wishes to raise $46,000 in cash and is considering two financing options: Clayton can sell $46,000 of bonds payable, or it can issue additional common stock for $46,000. To help in the decision process, Clayton’s management wants to determine the effects of each alternative on its current ratio and debt-to-assets ratio.

a-2. Compute the debt-to-assets ratio for Clayton’s management. (Round your answers to 1 decimal place.)

Currently:

If bonds are issued:

If stock is issued:

b. Assume that after the funds are invested, EBIT amounts to $12,500. Also assume the company pays $3,200 in dividends or $3,200 in interest depending on which source of financing is used. Based on a 40 percent tax rate, determine the amount of the increase in retained earnings that would result under each financing option.

Bonds:

Stock:
Answer #1

a-2: Debt to asset ratio:

Currently: (12000+56000)/(11000+88000) = 0.7

If bonds are issued: Total debt = 12000+56000+46000 = 114,000. Total assets = 11000+88000+46000 = 145,000. Thus ratio = 114,000/145,000 = 0.8

If stock is issued then ratio = (12000+56000)/145,000 = 0.5

b: Bonds: EBIT = 12,500. PBT = 12,500 – 3200 (interest) = 9300. Tax = 40% of 9300 = 3,720. Thus net income = 9300-3720 = 5,580. So retained earnings = $5,580

Stocks: EBIT = 12,500. As there is no interest PBT = 12,500 and tax = 40% of 12500 = 5000. So Net income = 12,500-5,000 = 7,500

Now retained earnings = Net income – dividends = 7500 – 3200 = 4,300. So retained earnings = $4,300